# Entanglement–magic separation in hybrid quantum circuits

### 04 Dec 2023 v1

@@ Abstract
Figure 1: Phase transitions in entanglement (a) compared to the possible phase transitions in entanglement and magic (b). Quantum circuits in phase I may lead to an advantage over classical computing, while circuits in phases II and III are amenable to simulation through tensor network methods, and circuits in phases III and IV can be tackled efficiently with the stabilizer formalism.
@@
The importance of entanglement has stimulated intense interest in its role not only in quantum information but also in other fields reaching from condensed matter to statistical mechanics and the quantum mechanics of black holes [13, 14, 15, 16, 17, 18, 19]. In some many body systems entanglement may display critical behavior even when all state–observables are featureless. One example of this are hybrid quantum circuits that consist of unitary gates interspersed with measurements [20, 21, 22, 23]. These circuits can exhibit so–called Measurement Induced Phase Transitions (MIPTs) in entanglement. There are multiple different circuits which show such MIPTs in entanglement, such as random Clifford [22, 23] or Haar–random circuits [21] interspersed with projective measurements. Both of these examples show the same qualitative behavior of entanglement, namely a volume law below a critical measurement rate and an area law above. However, from the perspective of a potential quantum advantage these two examples are very different. This is because even in the entanglement volume law phase the random Clifford circuit can be efficiently simulated classically using the stabilizer formalism, while the same is not true for the Haar–random circuit. Furthermore, MIPTs in entanglement have been related to the performance of quantum error correction codes, for which it has been found that hybrid circuits that show a volume law in entanglement encode quantum information, while circuits that show an area law destroy it irreversibly [24]. In this context, magic could further differentiate between quantum error correction codes that require non–Clifford resources (which are hard to implement in a fault–tolerant way) and codes that do not. It is thus natural to ask whether hybrid quantum circuits can also show MIPTs in magic, and if yes, how those phase transitions relate to the known MIPTs in entanglement.
@@
In this Letter we present a new kind of phase transition between a (sub)–extensive and area law scaling of magic in a hybrid quantum circuit controlled by the rate of measurement. Most notably, we find that this phase transition appears at a different critical measurement rate compared to the critical rate for the volume to area law transition of entanglement. This suggest that the mechanism that drives the observed magic phase transition is different from the mechanism driving the entanglement phase transition. From the perspective of a potential quantum advantage it is thus necessary to extend the picture of MIPTs in entanglement (see Fig. 1a) to MIPTs in entanglement and magic (see Fig. 1b). In particular, the hybrid quantum circuit studied in this Letter shows an entanglement phase transition from a (sub)–extensive to area law while the magic remains (sub)–extensive in both phases (phase I $\rightarrow$ phase II), as well as a magic phase transition from a (sub)–extensive to area law while the entanglement keeps an area law behavior (phase II $\rightarrow$ phase III).
@@
While phase transitions in magic have been reported previously in Refs. [25] and [26] we argue that the setup and the consequent phase transitions observed in this work are qualitatively different. The work by Leone et al. [25] studies to what extend Clifford circuits interspersed with T–gates can be transformed into circuits with all T–gates only in a small subset of qubits and does not involve any measurements. In Ref. [26] Niroula et al. study a random Clifford error correcting protocol with a stabilizer state as an input and a noise layer that injects magic through phase gates at a variable angle. This setup shows a phase transition of magic in the output state depending on the phase gate angle. Further analysis in Ref. [27] suggests that this phase transition stems from a phase transition in the success of error correction. Following from this, the transition presented in Ref. [26] could thus be witnessed through many different properties of a state, including magic, but also any other property, such as entanglement, would exhibit the same critical measurement rate set by the error correcting transition. Hence, to the best of our knowledge our work is the first to explicitly show the existence of an MIPT in magic that is independent of an MIPT in entanglement.
@@
where $\rho_{L}=\mathrm{tr}_{R}\left[|\psi\rangle\langle\psi|\right]$ is the reduced density matrix of the left half of the chain $L=\{1,\ldots,N/2-1\}$ with $\mathrm{tr}_{R}$ denoting the partial trace over the right half $R=\{N/2,\ldots,N\}$. While it is generally undisputed that the von–Neumann entropy is a good measure for bipartite entanglement in pure states, the choice of a “good measure” $\mathcal{M}$ for magic is under discussion in current literature [9, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. In this work we choose the recently introduced stabilizer entropy as a measure of magic [33].
@@
Figure 4: Phase diagram ($\eta$ vs. $p$) extracted from the numerical results for $\beta=1$ and $\eta=\{0.5,\ldots,8.0\}$. The bullets and crosses show the position of the entanglement and magic phase transition, respectively. For low measurement rates we find that both entanglement and magic follow a (sub)–extensive scaling (phase I), while for high measurement rates both follow an area law (phase III). However, we also find an extended region in between in which entanglement scales as an area law, but magic scales (sub)–extensively (phase II). The dotted line corresponds to the data shown in Fig. 3.
The data shown in Fig. 3 strongly suggests the existence of two distinct MIPTs. Figure 3a shows that in this log–linear plot for ¯$\bar{\mathcal{E}}_{\mathrm{ss}}(N)$, the line for $p=0.14$ tends upwards, while the line for $p=0.16$ appears approximately straight, and the lines for $p\geq 0.18$ bend downwards. This indicates a (sub)–extensive/log/area law in entanglement, respectively. This observation is consistent with the known measurement induced phase transition of entanglement for Clifford circuits, for which the addition of a vanishing density of T–gates appears not to significantly modify the critical measurement rate $p_{c}^{\mathrm{entgl.}}\simeq 0.16$ [22]. On the other hand, the values of magic for the same trajectories plotted in Fig. 3b also indicate (sub)–extensive/log/area law, but with a significantly higher critical measurement rate $p_{c}^{\mathrm{magic}}\simeq 0.22$. In particular, we note that for $p=0.18$ (marked with $\bigstar$) the results clearly indicate an area law for entanglement but a (sub)–extensive law for magic.
To extract the critical measurement rates from the data shown in Fig. 3 we fit each curve with two different scalings, corresponding to a (sub)–extensive and area law respectively and perform a so–called F–test [45]. For this, we fit $a$, $b$, and $\gamma$ in the function $f(N)=a+bN^{\gamma}$ with the constraint that $\gamma>0$ for the (sub)-extensive law and $\gamma\leq 0$ for the area law. For each fit we also compute the sum of the squared residuals per degree of freedom ($E=\chi^{2}/\mathsf{dof}$), which quantifies how well the fit matches the data. Hence for both entanglement and magic, all chosen values of $p$, and all chosen values of $\eta$ we have the quantities $E_{\mathrm{ext}}$ and $E_{\mathrm{area}}$ that quantify how well the data matches a (sub)–extensive and area law, respectively. Finally, using a linear interpolation between the discrete set of the $p$ values, we quote the value of $p$ for which both fits work equally well as the critical measurement rate.
@@
Acknowledgments.—We thank Titas Chanda, Mario Collura, Piotr Sierant, Poetri Tarabunga, and Xhek Turkeshi, for many helpful discussions. G. E. F. and R. F. acknowledge support from ERC under grant agreement n.101053159 (RAVE). M. D. and E. T. were partly supported by the MIUR Programme FARE (MEPH), by QUANTERA DYNAMITE PCI2022-132919, and by the EU-Flagship programme Pasquans2. M. D. and R. F. were partly supported by the PNRR MUR project PE0000023-NQSTI. M. D. and R. F. work was in part supported by the International Centre for Theoretical Sciences (ICTS) for participating in the program - Periodically and quasi-periodically driven complex systems (code: ICTS/pdcs2023/6).

### 11 Dec 2023 v2

@@ Abstract
Figure 1: Phase transitions in entanglement (a) compared to the possible phase transitions in entanglement and magic (b) in one–dimensional quantum circuits. Quantum circuits in phase I may lead to an advantage over classical computing, while circuits in phases II and III are amenable to simulation through tensor network methods, and circuits in phases III and IV can be tackled efficiently with the stabilizer formalism.
@@
The importance of entanglement has stimulated intense interest in its role not only in quantum information but also in other fields reaching from condensed matter to statistical mechanics and the quantum mechanics of black holes [13, 14, 15, 16, 17, 18, 19]. In some many body systems entanglement may display critical behavior even when all state–observables are featureless. One example of this are hybrid quantum circuits that consist of unitary gates interspersed with measurements [20, 21, 22, 23]. These circuits can exhibit so–called Measurement Induced Phase Transitions (MIPTs) in entanglement. There are multiple different circuits which show such MIPTs in entanglement, such as random Clifford [22, 23] or Haar–random circuits [21] interspersed with projective measurements. Both of these examples show the same qualitative behavior of entanglement, namely a volume law below a critical measurement rate and an area law above [24]. However, from the perspective of a potential quantum advantage these two examples are very different. This is because even in the entanglement volume law phase the random Clifford circuit can be efficiently simulated classically using the stabilizer formalism, while the same is not true for the Haar–random circuit. Furthermore, MIPTs in entanglement have been related to the performance of quantum error correction codes, for which it has been found that hybrid circuits that show a volume law in entanglement encode quantum information, while circuits that show an area law destroy it irreversibly [25]. In this context, magic could further differentiate between quantum error correction codes that require non–Clifford resources (which are hard to implement in a fault–tolerant way) and codes that do not. It is thus natural to ask whether hybrid quantum circuits can also show MIPTs in magic, and if yes, how those phase transitions relate to the known MIPTs in entanglement.
@@
In this Letter we present a new kind of phase transition between a (sub)–extensive and area law scaling of magic in a hybrid quantum circuit controlled by the rate of measurement. Most notably, we find that this phase transition appears at a different critical measurement rate compared to the critical rate for the volume to area law transition of entanglement. This suggest that the mechanism that drives the observed magic phase transition is different from the mechanism driving the entanglement phase transition. From the perspective of a potential quantum advantage it is thus necessary to extend the picture of MIPTs in entanglement (see Fig. 1a) to MIPTs in entanglement and magic (see Fig. 1b). In particular, the hybrid quantum circuit studied in this Letter shows an entanglement phase transition from a (sub)–extensive to area law while the magic remains (sub)–extensive in both phases [phase I $\rightarrow$ phase II], as well as a magic phase transition from a (sub)–extensive to area law while the entanglement keeps an area law behavior [phase II $\rightarrow$ phase III]. We note that in this paper we use the term “area law” for convenience to denote a constant scaling with system size, and that for magic there is no expectation of an according scaling in quantum circuits of higher dimensions.
@@
While phase transitions in magic have been reported previously in Refs. [26] and [27] we argue that the setup and the consequent phase transitions observed in this work are qualitatively different. The work by Leone et al. [26] studies to what extend Clifford circuits interspersed with T–gates can be transformed into circuits with all T–gates only in a small subset of qubits and does not involve any measurements. In Ref. [27] Niroula et al. study a random Clifford error correcting protocol with a stabilizer state as an input and a noise layer that injects magic through phase gates at a variable angle. This setup shows a phase transition of magic in the output state depending on the phase gate angle. Further analysis reported by Turkeshi and Sierant [28] suggests that this phase transition stems from a phase transition in the success of error correction. Following from this, the transition presented in Ref. [27] could thus be witnessed through many different properties of a state, including magic, but also any other property, such as entanglement, would exhibit the same critical measurement rate set by the error correcting transition. Hence, to the best of our knowledge our work is the first to explicitly show the existence of an MIPT in magic that is independent of an MIPT in entanglement.
@@
where $\rho_{L}=\mathrm{tr}_{R}\left[|\psi\rangle\langle\psi|\right]$ is the reduced density matrix of the left half of the chain $L=\{1,\ldots,N/2-1\}$ with $\mathrm{tr}_{R}$ denoting the partial trace over the right half $R=\{N/2,\ldots,N\}$. While it is generally undisputed that the von–Neumann entropy is a good measure for bipartite entanglement in pure states, the choice of a “good measure” $\mathcal{M}$ for magic is under discussion in current literature [9, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. In this work we choose the recently introduced stabilizer entropy as a measure of magic [36].
@@
Figure 4: Phase diagram ($\eta$ vs. $p$) extracted from the numerical results for $\beta=1$ and $\eta=\{0.5,\ldots,8.0\}$. The bullets and crosses show the position of the entanglement and magic phase transition, respectively. We present the details of the numerical error estimation (displayed as error bars) in the SM [48]. The colored regions resulting from this data correspond to the phases I, II, and III shown in Fig. 1. For low measurement rates we find that both entanglement and magic follow a (sub)–extensive scaling (phase I), while for high measurement rates both follow an area law (phase III). However, we also find an extended region in between in which entanglement scales as an area law, but magic scales (sub)–extensively (phase II). The dotted line corresponds to the data shown in Fig. 3.
@@
The data shown in Fig. 3 strongly suggests the existence of two distinct MIPTs. Figure 3a shows that in this log–linear plot for ¯$\bar{\mathcal{E}}_{\mathrm{ss}}(N)$, the line for $p=0.14$ tends upwards, while the line for $p=0.16$ appears approximately straight, and the lines for $p\geq 0.18$ bend downwards. This indicates a (sub)–extensive/log/area law in entanglement, respectively. This observation is consistent with the known measurement induced phase transition of entanglement for Clifford circuits, for which the addition of a vanishing density of T–gates appears not to significantly modify the critical measurement rate $p_{c}^{\mathrm{entgl.}}=0.15995(10)$ [22, 51]. On the other hand, the values of magic for the same trajectories plotted in Fig. 3b also indicate (sub)–extensive/log/area law, but with a significantly higher critical measurement rate $p_{c}^{\mathrm{magic}}\simeq 0.22$. In particular, we note that for $p=0.18$ (marked with $\bigstar$) the results clearly indicate an area law for entanglement but a (sub)–extensive law for magic.
As the scaling of magic near the critical point is yet unknown we do not perform any scaling collapse in this work. We instead focus on the extraction of the critical measurement rates from the data shown in Fig. 3 by fitting each curve with two different functions, corresponding to a (sub)–extensive and area law respectively and perform a so–called F–test [52]. For this, we fit $a$, $b$, and $\gamma$ in the function $f(N)=a+bN^{\gamma}$ with the constraint that $\gamma>0$ for the (sub)-extensive law and $\gamma\leq 0$ for the area law. For each fit we also compute the sum of the squared residuals per degree of freedom ($E=\chi^{2}/\mathsf{dof}$), which quantifies how well the fit matches the data. Hence for both entanglement and magic, all chosen values of $p$, and all chosen values of $\eta$ we have the quantities $E_{\mathrm{ext}}$ and $E_{\mathrm{area}}$ that quantify how well the data matches a (sub)–extensive and area law, respectively. Finally, using a linear interpolation between the discrete set of the $p$ values, we quote the value of $p$ for which both fits work equally well as the critical measurement rate.
@@
Acknowledgments.—We thank Lorenzo Piroli, Xhek Turkeshi, and Christopher White for helpful comments on the manuscript. Also, we thank Titas Chanda, Mario Collura, Piotr Sierant, Poetri Tarabunga, and Xhek Turkeshi, for many insightful discussions. G. E. F. and R. F. acknowledge support from ERC under grant agreement n.101053159 (RAVE). M. D. and E. T. were partly supported by the MIUR Programme FARE (MEPH), by QUANTERA DYNAMITE PCI2022-132919, and by the EU-Flagship programme Pasquans2. M. D. and R. F. were partly supported by the PNRR MUR project PE0000023-NQSTI. M. D. and R. F. work was in part supported by the International Centre for Theoretical Sciences (ICTS) for participating in the program - Periodically and quasi-periodically driven complex systems (code: ICTS/pdcs2023/6).
Note added: After uploading this manuscript to arXiv, we became aware of a related work by Bejan, McLauchlan and Beri [53] that also discusses the interplay of magic and entanglement transitions in random unitary circuits. While the details of the works largely differ, they both observe that entanglement and magic transitions are in general distinct.